OpenStudy (anonymous):

ok another one that has me stumped and I am sure it is in the operations -9-2=6x+2-5x x=

OpenStudy (anonymous):

-11=x+2 x=-13.

OpenStudy (anonymous):

Ok but I am trying to understand the order of operation and and how to come up with that answer... THis is so confusing to me

OpenStudy (princeharryyy):

take all the constants on the one side and the variables on the other and solve @michelebeme2

OpenStudy (anonymous):

Combine like terms. (-9-2) = 11 on the left. (6x-5x) = x and the constant 2 on the right. Then just subtract two from the left and right to combine the like terms and you get x = -13.

OpenStudy (anonymous):

ok so it will always be just x on the left then ? You dont add 6x + 5x to get 11x its just x ?

OpenStudy (princeharryyy):

its a - sign in between not a + one @michelebeme2

OpenStudy (anonymous):

ok but if it is a - sign then wouldn't you subtract 6x - 5x to get 1 x ? oh wait I got it one x or x is the same thing?

OpenStudy (princeharryyy):

yes they are @michelebeme2